OpenStudy (anonymous):

a person's mass while dieting decreases by 5%. Exercise creates muscle and reduces fat, causing the person's density to increase by 2%. calculate the percent change in the persons volume.

OpenStudy (anonymous):

density = mass/volume volume = mass/density let v1 = original, v2 = modified v2/v1 = (m2/m1)/(d2/d1) = (0.95)/(1.02) = 0.93 so volume goes down 7% seems odd, though

OpenStudy (anonymous):

thanks ! :))

OpenStudy (anonymous):

You are very welcome.